OpenStudy (anonymous):

the graph of f(x) = 2^(x + 4) shifts five units to the right when it is replaced with the graph of f(x) = 2^(x - k). What is the value of k?

OpenStudy (anonymous):

@jim_thompson5910 can you help me with this one too? Its -5 right?

jimthompson5910 (jim_thompson5910):

hint: to shift f(x) to the right 5 units, you replace x with x-5 so we need to find f(x-5)

OpenStudy (anonymous):

Okay, \[f(x - 5) = 2^{ -5 +4}\]?

jimthompson5910 (jim_thompson5910):

where did x go on the right side?

OpenStudy (anonymous):

Wait a second.. I replaced it with -5, and I shouldn't have.

OpenStudy (anonymous):

\[f(x -5) = 2^{x -5 + 4}\]

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

now simplify

OpenStudy (anonymous):

\[-5+4 = -1\]\[f(x - 5) = 2^{x -1}\] and I'm drawing a blank here.

OpenStudy (anonymous):

could I add 5 on both sides?

jimthompson5910 (jim_thompson5910):

compare 2^(x-1) to 2^(x-k)

OpenStudy (anonymous):

Does that mean that the answer is 1?

jimthompson5910 (jim_thompson5910):

it does

jimthompson5910 (jim_thompson5910):

k = 1

OpenStudy (anonymous):

My initial answer moved 9 units, not 5.

jimthompson5910 (jim_thompson5910):

yeah a graph confirms it with k = 5, you move over 9 units to the right

OpenStudy (anonymous):

Thank you, I really appreciate your help. You're like the only person I understand when I'm on here.

jimthompson5910 (jim_thompson5910):

you're welcome